OpenStudy (anonymous):

ANSWERS GET MEDALS What is the equation of a circle formed by the set of all points 9 units away from the point (6, 2)?

OpenStudy (amistre64):

well the radius is 2; and the center is 6 and 2 :)

OpenStudy (amistre64):

(x-Cx)^2 + (y-Cy)^2 = r^2

OpenStudy (amistre64):

radius is 9 lol ..... im gonna poke my brain with a qtip for that

OpenStudy (anonymous):

Yeah, I was just about to ask about that..

OpenStudy (anonymous):

(x-6)^2+(y-2)^2 = 81